A person does 360J of work in 3.2s to accelerate a bike at 13m/s^2. This changed the velocity by 31m/s. How heavy is the bike. The bike moved 26m


1
Expert's answer
2022-06-06T08:52:59-0400
"W=Fd"

"F=\frac{W}{d}=\frac{360}{26}=13.8\:\rm N"

"F=ma"

"m=\frac{F}{a}=\frac{13.8}{13}=1.1\:\rm kg"